OpenStudy (anonymous):

Solving Special Systems Classify system. Give the number of solutions. y=3(x-1) -y+3x=3

OpenStudy (texaschic101):

y = 3(x - 1) y = 3x - 3 -y + 3x = 3 -y = -3x + 3 -- multiply by -1 y = 3x - 3 same line...infinitely many solutions